OpenStudy (anonymous):

from the equation of a circle (x-4)^2+(y-8)^2=64, which are the center and the radius?

OpenStudy (anonymous):

centre (4,8) and radius=sqrt(640=8 units

OpenStudy (anonymous):

(x-a)^2+(y-b)^2=r^2 (a,b) is the center r is the radius

OpenStudy (anonymous):

for centre (h,k) and raius r the eq is (x-h)^2 +(y-k)^2=r^2

OpenStudy (anonymous):

hrmm... what's the answer? o.o

OpenStudy (agent0smith):

Compare your equation: \[(x-4)^{2}+(y-8)^{2} = 64\] to the standard form: \[(x-h)^{2} +(y-k)^{2}=r ^{2}\] Where, as matricked and wacsam said, the center is (h, k) and the radius is r.